# latex‘!h’ float specifier changed to ‘!ht’

### v1

@@ 5 Real-data analysis
We validate our framework on 6 real-world examples. More specifically, we used the publicly available regression datasets used as validation by [16] and [14]. These datasets include 4 spatial datasets on election results, income, unemployment as well as education levels across the counties in the US. These datasets also include 2 transportation networks, based on traffic data in the cities of Anaheim and Chicago. Each vertex represents a directed lane, and two lanes that meet at the same intersection are connected. In these datasets, the signal $Y$ represents the traffic flow in the lane. Finally, the Twitch dataset captures friendships amongst Twitch streamers in Portugal, while the node signal corresponds to the logarithm of the number of viewers for each streamer. A table listing the datasets and their properties can be found in the Appendix, highlighting the different topological characteristics of these datasets. In particular, spatial datasets tend to have a regular degree, while the degree distribution is highly skewed in the twitch datasets.
@@
We use our estimator in the prediction setting, where the value at the node of interest is unobserved. To this end, we select at random a set of 500 nodes for training (i.e. selection of the most appropriate value of neighborhood size $L$) and 500 for validation. Figure 4 shows the MSE error as a function of the neighborhood size. Interestingly, we note that the optimal neighborhood size is reached at $L=2$ for all spatial graphs. In these settings, the degree distribution is relatively homogeneous across the nodes (see statistics in the Appendix), and the convolutions $S$ and $T$ give reasonably similar results. Interestingly, however, in the social network example, the convolution yields drastically different effects.
@@
We also test our estimator in the denoising setting. To avoid data snooping, we replace the value at the selected nodes with one of their direct neighbours. In the homophilic graph setting, this is indeed akin to assuming that direct neighbours have roughly the same distribution of the node of interest. Therefore, replacing the node by one of its neighbors ensures that we are generating an independent copy of the graph with the same distribution. The results are presented in the Appendix.
@@
Figure 4: MSE as a function of the neighborhood size. The nodes denote the mean MSE over 50 random splits of the data into training and validation sets, and the error bars denote interquartile ranges. Note that the $y$-axis does not start at 0.
6 Discussion
In this work we pave the way towards a statistical foundation of graph convolutional networks (GCNs). Graph convolutional networks perform weighted local averaging over neighboring nodes where the neighborhood size is controlled by the depth of the GCN. Studying node regression, we showed that the variance of the GCN is a weighted sum over paths. This link allows to derive lower and upper bounds of the variance based on local graph properties. Our analysis reveals some non-robustness of GCNs in the sense that small perturbations of these graph properties can lead to huge changes in the variance.
Acknowledgements
@@
The research of J. S.-H. and J. C. was supported by the NWO Vidi grant VI.Vidi.192.021. The research of C. D. was funded by the National Science Foundation (Award Number 2238616), as well as the resources provided by the University of Chicago’s Research Computing Center. The work of O. K. was funded by CY Initiative (grant “Investissements d’Avenir” ANR-16-IDEX0008) and Labex MME-DII(ANR11-LBX-0023-01).
@@ A.1 Proof of Theorem 1
Recall that as the noise ${\varepsilon}_{i}$ is centered, with inequality (7), we have for any fixed parameter $W\in{\mathbb{R}}$ and matrix $P\in{\mathbb{R}}^{n\times n}$,
${\mathbb{E}}\left[{\|WP\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]\leq\Big{(}W\big{\|}P\mathbf{f^{*}}-\mathbf{f^{*}}\big{\|}_{2}+|1-W|\cdot\|\mathbf{f^{*}}\|_{2}\Big{)}^{2}+W^{2}{\mathbb{E}}\left[{\|P{\bm{{\varepsilon}}}\|_{2}^{2}}\right].$ (1)
We first consider the case $P=S^{L}$ for $L=1,2,\ldots$. To further bound (1) for $P=S^{L}$, we now prove by induction on $L$ that
$\|S^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\|_{\infty}\leq L\Delta,$ (2)
@@
$\tilde{A}=\begin{pmatrix}a_{11}&a_{12}&\ldots&a_{1n}\\ a_{21}&a_{22}&\ldots&a_{2n}\\ \ldots&\ldots&\ldots&\ldots\\ a_{n1}&a_{n2}&\ldots&a_{nn}\\ \end{pmatrix},$ (3)
@@
$S=\tilde{D}^{-1}\tilde{A}=\begin{pmatrix}\frac{a_{11}}{d_{1}+1}&\frac{a_{12}}{d_{1}+1}&\ldots&\frac{a_{1n}}{d_{1}+1}\\ \frac{a_{21}}{d_{2}+1}&\frac{a_{22}}{d_{2}+1}&\ldots&\frac{a_{2n}}{d_{2}+1}\\ \ldots&\ldots&\ldots&\ldots\\ \frac{a_{n1}}{d_{n}+1}&\frac{a_{n2}}{d_{n}+1}&\ldots&\frac{a_{nn}}{d_{n}+1}\\ \end{pmatrix},$ (4)
@@
$\|S^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\|_{2}\leq\sqrt{n}\|S^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\|_{\infty}\leq\sqrt{n}L\Delta.$ (5)
@@
${\mathbb{E}}\left[{\|S^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]=\mathop{\rm Tr}\nolimits((S^{L})^{\top}S^{L})=\|S^{L}\|_{F}^{2}.$ (6)
Plugging (5) and (6) into (1) with $P=S^{L}$, we deduce that
$\frac{1}{n}{\mathbb{E}}\left[{\|WS^{L}\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]$ $\leq\frac{1}{n}\Big{(}W\big{\|}S^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\big{\|}_{2}+|1-W|\cdot\|\mathbf{f^{*}}\|_{2}\Big{)}^{2}+\frac{W^{2}}{n}{\mathbb{E}}\left[{\|S^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]$
$\leq\left(WL\Delta+\frac{|1-W|}{\sqrt{n}}\cdot\|\mathbf{f^{*}}\|_{2}\right)^{2}+W^{2}\frac{\|S^{L}\|_{F}^{2}}{n},$
which completes the proof of (10).
@@
$T^{L}\mathbf{f^{*}}-\mathbf{f^{*}}$ $=\tilde{D}^{1/2}S^{L}\tilde{D}^{-1/2}\mathbf{f^{*}}-\mathbf{f^{*}}=\tilde{D}^{1/2}\big{(}S^{L}-I_{n}\big{)}\tilde{D}^{-1/2}\mathbf{f^{*}}.$ (7)
Since $\mathbf{f^{*}}$ satisfies (9) in Assumption 1, the vector $\tilde{D}^{-1/2}f^{*}$ satisfies (8) in Assumption 1. Using (2) and (7),
@@
$=\sqrt{2|\mathcal{E}|+n}\cdot L\Delta,$ (8)
@@
For the variance term, we can proceed similarly as in (6) and obtain
${\mathbb{E}}\left[{\|T^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]=\mathop{\rm Tr}\nolimits((T^{L})^{\top}T^{L})=\|T^{L}\|_{F}^{2}.$ (9)
Plugging (8) and (9) into (1) with $P=T^{L}$, we deduce
$\frac{1}{n}{\mathbb{E}}\left[{\|WT^{L}\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]$ $\leq\frac{1}{n}\Big{(}W\big{\|}T^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\big{\|}_{2}+|1-W|\cdot\|\mathbf{f^{*}}\|_{2}\Big{)}^{2}+\frac{W^{2}}{n}{\mathbb{E}}\left[{\|T^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]$
$\leq\left(WL\Delta\sqrt{\frac{2|\mathcal{E}|}{n}+1}+\frac{|1-W|}{\sqrt{n}}\cdot\|\mathbf{f^{*}}\|_{2}\right)^{2}+W^{2}\frac{\|T^{L}\|_{F}^{2}}{n},$
proving (11).
@@ A.2 Proof of Theorem 2
$(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j),$ (10)
which connects node $i$ to node $j$ through nodes $\ell_{1},\ldots,\ell_{L-1}\in\mathcal{V}$ and has a length of $L$. The path in (10) also implies that $\ell_{1}\in{\mathcal{N}}^{1}(i),\ell_{2}\in{\mathcal{N}}^{1}(\ell_{1}),\ldots,j\in{\mathcal{N}}^{1}(\ell_{L-1})$. Therefore, if condition (8) in Assumption 1 holds, we have, for any $j\in{\mathcal{N}}^{L}(i)$,
$|f_{i}^{*}-f_{j}^{*}|\leq|f_{i}^{*}-f_{\ell_{1}}^{*}|+|f_{\ell_{1}}^{*}-f_{\ell_{2}}^{*}|+\cdots+|f_{\ell_{L-1}}^{*}-f_{j}^{*}|\leq L\Delta.$ (11)
Since the noise variables ${\varepsilon}_{i}$ are assumed to be uncorrelated, centered, and have a variance of one, we can derive using (11) that for any $i=1,\ldots,n$,
@@
$\leq L^{2}\Delta^{2}+\frac{1}{|{\mathcal{N}}^{L}(i)|}.$ (12)
Summing (12) for all $i=1,\ldots,n$ finally yields
@@ B.1 Associated weights for $S^{L}$
Recall that we can represent $S$ as shown in (4), where $a_{ij}$ for $i\not=j$ indicates whether node $i$ is connected to node $j$ by an edge in $\mathcal{E}$, and $a_{ii}=1$ since self-loops are included. Similarly, we define $a_{(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j)}:=1$, if there exists a path with self-loops of length $L$ from node $i$ to node $j$ via nodes $\ell_{1},\ldots\ell_{L-1}\in\{1,\ldots,n\}$; otherwise $a_{(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j)}:=0$. In particular, when $L=1$, we have $a_{ij}=a_{(i\to j)}$.
@@
$s^{L}_{ij}=\sum_{\ell_{1},\ldots,\ell_{L-1}=1}^{n}\frac{a_{(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j)}}{(d_{i}+1)(d_{\ell_{1}}+1)\cdot\ldots\cdot(d_{\ell_{L-1}}+1)},$ (13)
where $d_{i},d_{\ell_{1}},\ldots,d_{\ell_{L-1}}$ denote the edge degrees of nodes $i,\ell_{1},\ldots,\ell_{L-1}$, respectively. The equality (13) then implies that the associated weight for each path is given by
@@
One can observe from (4) that the conclusion holds for $L=1$. For the induction step, assume the claim holds until $L-1$, in the sense that, for any $11\leq k\leq L-$ 1, and $i,j=1,\ldots,n$,

### v2

@@ 6 REAL-DATA ANALYSIS
We validate our framework on 6 real-world examples. More specifically, we used the publicly available regression datasets used as validation by Jia and Benson, (2020) and Huang et al., (2024). These datasets include 3 spatial datasets on election results, income as well as education levels across the counties in the United States. We also include 2 transportation networks, based on traffic data in the cities of Anaheim and Chicago. Each vertex represents a directed lane, and two lanes that meet at the same intersection are connected. In these datasets, the response $Y_{i}$ represents the traffic flow in the lane. Finally, the Twitch dataset captures friendships among Twitch streamers in Portugal, while the node signal corresponds to the logarithm of the number of viewers for each streamer. Figure 5 provides details on the datasets and their properties, emphasizing their distinct topological characteristics. In particular, spatial datasets typically have a regular degree distribution, whereas the Twitch dataset exhibits a highly skewed degree distribution.
@@
We use our estimator in the prediction setting, where the value at the node of interest is unobserved. In this setting, we predict $\hat{Y}_{i}$ using different convolutions: $\hat{Y}_{i}=\sum_{j\in\mathcal{N}(i)}w_{ij}Y_{j}$, where the weights in the sum depend on the convolution. To emulate a real-life training scenario, we randomly select a set of 500 nodes for training and another 500 for validation. Figure 4 shows the mean squared error on the validation set as a function of the neighborhood size. Interestingly, we observe that the optimal neighborhood size is reached at $L=2$ for all spatial graphs. In these cases, the degree distribution is relatively homogeneous across nodes (see statistics in the appendix), and the convolutions $S$ and $T$ produce fairly similar results. However, in the social network example, different aggregation operators lead to drastically different effects, with the GraphSAGE convolution outperforming the other two.
@@
We also test our estimator as a denoiser. In this setting, we denoise the observed value $Y_{i}$ as $\hat{Y}_{i}=\sum_{j\in\mathcal{N}^{1}(i)}w_{ij}Y_{j}$, with the weights depending on the convolution. To avoid data snooping between the training and test sets, we replace the values of the test nodes with those of one of their direct neighbors. Assuming that direct neighbors have roughly the same distribution, this modification ensures that no information from the training set permeates the test set while creating another version of the same graph. The results are presented in the appendix in Figure  6.
7 CONCLUSION
In this work, we lay the groundwork for a statistical understanding of GCNs. GCNs perform weighted local averaging over neighboring nodes, with the neighborhood size determined by the depth $L$ of the network. In the context of denoising, we characterize a bias-variance type trade-off risk bound in terms of $L$ for both the original GCN and GraphSAGE convolution. We showed that the variance term of the GraphSAGE convolution is lower-bounded by its counterpart from the local averaging estimator. We propose investigating the variance term from a new perspective: walk analysis, which enables the derivation of both lower and upper bounds of the variance based on local graph topologies. Our analysis reveals a lack of robustness in GCNs, in that small perturbations in the graph structure can lead to substantial changes in the variance.
Figure 1: Optimal number of convolutions as a function of the roughness $\|\Delta{\bf f^{*}}\|_{2}=\sqrt{[\sum_{(i,j)\in\mathcal{E}}(f_{i}^{*}-f_{j}^{*})^{2}]/|\mathcal{E}|}$ of $\mathbf{f}^{*}$ on the latent variable graph with $\sigma^{2}=2$. Each subplot corresponds to a different sparsification level, with 0 representing the original graph and 0.75 representing a graph with 75% of its edges removed.
Figure 2: The variance decay at the root of the tree as a function of $L$ under the signal $f_{i}^{*}=2\cos(U_{i\cdot}\beta)$, with $\beta=(-1,1)^{\top}$, colored by degree value. The $y$-axis is shared across all 3 plots.
Figure 3: The variance decay at the root of the tree as a function of $L$ under the signal $f_{i}^{*}=2\cos(U_{i\cdot}\beta)$, with $\beta=(-1,1)^{\top}$, and its behavior after adding cycles.
@@ 7 CONCLUSION
Figure 4: Mean Squared Error (MSE) as a function of the neighborhood size. The nodes denote the mean MSE over 50 random splits of the data into training and validation sets, and the error bars denote interquartile ranges.
Acknowledgments
@@
We thank the anonymous reviewers for their helpful comments. The research of J. S.-H. and J. C. was supported by the NWO Vidi grant VI.Vidi.192.021. The research of C. D. was funded by the National Science Foundation (Award Number 2238616), as well as the resources provided by the University of Chicago’s Research Computing Center. The work of O. K. was funded by CY Initiative (grant “Investissements d’Avenir” ANR-16-IDEX0008) and Labex MME-DII(ANR11-LBX-0023-01). Part of this work was carried out while J. S.-H. and J. C. were visiting the Simons Institute for the Theory of Computing in Berkeley.
@@ A.1 Proof of Theorem 1
Recall that as the noise ${\varepsilon}_{i}$ is centered, with inequality (8) in the paper, we have for any fixed parameter $W\in{\mathbb{R}}$ and matrix $P\in{\mathbb{R}}^{n\times n}$,
${\mathbb{E}}\left[{\|WP\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]\leq\Big{(}|W|\cdot\big{\|}P\mathbf{f^{*}}-\mathbf{f^{*}}\big{\|}_{2}+|1-W|\cdot\|\mathbf{f^{*}}\|_{2}\Big{)}^{2}+W^{2}{\mathbb{E}}\left[{\|P{\bm{{\varepsilon}}}\|_{2}^{2}}\right].$ (A.1)
We first consider the case $P=S^{L}$ for $L=1,2,\ldots$. To further bound (A.1) for $P=S^{L}$, we now prove by induction on $L$ that
$\|S^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\|_{\infty}\leq L\Delta,$ (A.2)
@@
$\tilde{A}=\begin{pmatrix}a_{11}&a_{12}&\ldots&a_{1n}\\ a_{21}&a_{22}&\ldots&a_{2n}\\ \ldots&\ldots&\ldots&\ldots\\ a_{n1}&a_{n2}&\ldots&a_{nn}\\ \end{pmatrix},$ (A.3)
@@
$S=\tilde{D}^{-1}\tilde{A}=\begin{pmatrix}\frac{a_{11}}{d_{1}+1}&\frac{a_{12}}{d_{1}+1}&\ldots&\frac{a_{1n}}{d_{1}+1}\\ \frac{a_{21}}{d_{2}+1}&\frac{a_{22}}{d_{2}+1}&\ldots&\frac{a_{2n}}{d_{2}+1}\\ \ldots&\ldots&\ldots&\ldots\\ \frac{a_{n1}}{d_{n}+1}&\frac{a_{n2}}{d_{n}+1}&\ldots&\frac{a_{nn}}{d_{n}+1}\\ \end{pmatrix},$ (A.4)
@@
$\|S^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\|_{2}\leq\sqrt{n}\|S^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\|_{\infty}\leq\sqrt{n}L\Delta.$ (A.5)
@@
${\mathbb{E}}\left[{\|S^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]=\mathop{\rm Tr}\nolimits((S^{L})^{\top}S^{L})=\|S^{L}\|_{F}^{2}.$ (A.6)
Plugging (A.5) and (A.6) into (A.1) with $P=S^{L}$, we deduce that
$\frac{1}{n}{\mathbb{E}}\left[{\|WS^{L}\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]$ $\leq\frac{1}{n}\Big{(}|W|\cdot\big{\|}S^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\big{\|}_{2}+|1-W|\cdot\|\mathbf{f^{*}}\|_{2}\Big{)}^{2}+\frac{W^{2}}{n}{\mathbb{E}}\left[{\|S^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]$
$\leq\left(|W|\cdot L\Delta+\frac{|1-W|}{\sqrt{n}}\cdot\|\mathbf{f^{*}}\|_{2}\right)^{2}+W^{2}\frac{\|S^{L}\|_{F}^{2}}{n}$
$=\operatorname{Bias}_{S}^{2}(L,W)+W^{2}\frac{\|S^{L}\|_{F}^{2}}{n},$
which concludes the proof of the first upper bound in Theorem 1 of the paper.
@@
$T^{L}\mathbf{f^{*}}-\mathbf{f^{*}}$ $=\tilde{D}^{1/2}S^{L}\tilde{D}^{-1/2}\mathbf{f^{*}}-\mathbf{f^{*}}=\tilde{D}^{1/2}\big{(}S^{L}-I_{n}\big{)}\tilde{D}^{-1/2}\mathbf{f^{*}}.$ (A.7)
Since $\mathbf{f^{*}}$ satisfies (10) in Assumption 1, the vector $\tilde{D}^{-1/2}f^{*}$ satisfies (9) in Assumption 1. Using (A.2) and (A.7),
@@
$=\sqrt{2|\mathcal{E}|+n}\cdot L\Delta,$ (A.8)
@@
For the variance term, we can proceed similarly as in (A.6) and obtain
${\mathbb{E}}\left[{\|T^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]=\mathop{\rm Tr}\nolimits((T^{L})^{\top}T^{L})=\|T^{L}\|_{F}^{2}.$ (A.9)
Plugging (A.8) and (A.9) into (A.1) with $P=T^{L}$, we deduce
$\frac{1}{n}{\mathbb{E}}\left[{\|WT^{L}\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]$ $\leq\frac{1}{n}\Big{(}|W|\cdot\big{\|}T^{L}\mathbf{f^{*}}-\mathbf{f^{*}}\big{\|}_{2}+|1-W|\cdot\|\mathbf{f^{*}}\|_{2}\Big{)}^{2}+\frac{W^{2}}{n}{\mathbb{E}}\left[{\|T^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]$
$\leq\left(|W|\cdot L\Delta\sqrt{\frac{2|\mathcal{E}|}{n}+1}+\frac{|1-W|}{\sqrt{n}}\cdot\|\mathbf{f^{*}}\|_{2}\right)^{2}+W^{2}\frac{\|T^{L}\|_{F}^{2}}{n}$
$=\operatorname{Bias}_{T}^{2}(L,W)+W^{2}\frac{\|T^{L}\|_{F}^{2}}{n},$
proving the second upper bound in Theorem 1.
@@ A.2 Proof of Theorem 2
$(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j),$ (A.10)
which connects node $i$ to node $j$ through nodes $\ell_{1},\ldots,\ell_{L-1}\in\mathcal{V}$ and has a length of $L$. The path in (A.10) also implies that $\ell_{1}\in{\mathcal{N}}^{1}(i),\ell_{2}\in{\mathcal{N}}^{1}(\ell_{1}),\ldots,j\in{\mathcal{N}}^{1}(\ell_{L-1})$. Therefore, if condition (9) in Assumption 1 holds, we have, for any $j\in{\mathcal{N}}^{L}(i)$,
$|f_{i}^{*}-f_{j}^{*}|\leq|f_{i}^{*}-f_{\ell_{1}}^{*}|+|f_{\ell_{1}}^{*}-f_{\ell_{2}}^{*}|+\cdots+|f_{\ell_{L-1}}^{*}-f_{j}^{*}|\leq L\Delta.$ (A.11)
Since the noise variables ${\varepsilon}_{i}$ are assumed to be uncorrelated, centered, and have a variance of one, we can derive using (A.11) that for any $i=1,\ldots,n$,
@@
$\leq L^{2}\Delta^{2}+\frac{1}{|{\mathcal{N}}^{L}(i)|}.$ (A.12)
Summing (A.12) for all $i=1,\ldots,n$ finally yields
@@ B.1 Associated Weights for $S^{L}$
Recall that we can represent $S$ as shown in (A.4), where $a_{ij}$ for $i\not=j$ indicates whether node $i$ is connected to node $j$ by an edge in $\mathcal{E}$, and $a_{ii}=1$ since self-loops are included. Similarly, we define $a_{(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j)}:=1$, if there exists a path with self-loops of length $L$ from node $i$ to node $j$ via nodes $\ell_{1},\ldots\ell_{L-1}\in\{1,\ldots,n\}$; otherwise $a_{(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j)}:=0$. In particular, when $L=1$, we have $a_{ij}=a_{(i\to j)}$.
@@
$s^{L}_{ij}=\sum_{\ell_{1},\ldots,\ell_{L-1}=1}^{n}\frac{a_{(i\to\ell_{1}\to\ldots\to\ell_{L-1}\to j)}}{(d_{i}+1)(d_{\ell_{1}}+1)\cdot\ldots\cdot(d_{\ell_{L-1}}+1)},$ (A.13)
where $d_{i},d_{\ell_{1}},\ldots,d_{\ell_{L-1}}$ denote the edge degrees of nodes $i,\ell_{1},\ldots,\ell_{L-1}$, respectively. The equality (A.13) then implies that the associated weight for each path is given by
@@
One can observe from (A.4) that the conclusion holds for $L=1$. For the induction step, assume the claim holds until $L-1$, in the sense that, for any $11\leq k\leq L-$ 1, and $i,j=1,\ldots,n$,